OpenStudy (anonymous):

simplify. please show all of your work. 5c - 3bc + b - (-c -b - 3bc)

OpenStudy (bahrom7893):

5c - 3bc + b + c + b + 3bc = 5c + c + b + b - 3bc + 3bc = 6c + 2b + 0 = 6c + 2b = 2b+6c